OpenStudy (ibbutibbu.):

Can I please get some help? The following function defines a recursive sequence. f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? A) -4, 12, -28, 60, … B) -4, -12, -28, -60, … C) -4, 12, -18, 54, … D) -4, 12, -18, -54, …

OpenStudy (anonymous):

A

OpenStudy (ibbutibbu.):

Thank You!

OpenStudy (ibbutibbu.):

I have one more question please

OpenStudy (ibbutibbu.):

The graph below shows the distance (y) between Hannah and her home, in km, after a certain time (x), in hours. Which of the following statements best describes Hannah's distance from her home? It is decreasing in the interval 2 < x < 3 hours. It is increasing in the interval 2 < x < 3 hours. It is increasing in the interval 2 < x < 4 hours. It is decreasing in the interval 2 < x < 4 hours.

OpenStudy (anonymous):

A

OpenStudy (ibbutibbu.):

Thank You!

OpenStudy (ibbutibbu.):

The first four terms of a sequence are shown below. 7, 4, 1, -2 Which of the following functions best defines this sequence? (4 points) f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1

OpenStudy (anonymous):

C

OpenStudy (ibbutibbu.):

ty

OpenStudy (ibbutibbu.):

William raced up a flight of stairs at the rate of 3 stairs per second. The function below shows the total number of stairs f(t) that he is able to climb in t seconds. f(t) = 3t + 4 Which statement best describes the y intercept of the graph of the function? (3 points) It is equal to 3 and represents the number of stairs William had already climbed before he started racing up. It is equal to 3 and represents the number of stairs William can climb in 1 second. It is equal to 4 and represents the number of stairs William had already climbed before he started racing up. It is equal to 7 and represents the number of stairs William can climb in 1 second.

OpenStudy (anonymous):

C

OpenStudy (ibbutibbu.):

lol thanks